OpenStudy (anonymous):

Show that if n is a positive integer then Question below:

OpenStudy (anonymous):

\[(\sum_{d|n}^{} \tau(d) )^2 = \sum_{d|n}^{} \tau(d)^3 \]

OpenStudy (blurbendy):

i'm unfamiliar with the notation: d|n care to elaborate?

OpenStudy (blurbendy):

divides?

OpenStudy (anonymous):

yea its the divisor

OpenStudy (anonymous):

so taking the summatory function of tau then sqauring it is equal to the summatory of tau cubed

OpenStudy (blurbendy):

im not sure i understand how the whole expression on the left can be squared. did you just mean to square tau(d)^2 ?

OpenStudy (anonymous):

nope its written correctly the lhs is completely squared

OpenStudy (blurbendy):

interesting..

OpenStudy (anonymous):

this is supposed to be a hard problem so its gonna take a while to figure out lol

OpenStudy (blurbendy):

if it wasn't so late, i'd have more enthusiasm ;) im a little fuzzy right now. ill come back to this tomorrow if you dont already have an answer by then

OpenStudy (anonymous):

oh thats alright thanks for looking/trying at the problem ^_^

OpenStudy (kc_kennylau):

what is the tau function? :O

OpenStudy (anonymous):

the tau function is number of divisors

OpenStudy (anonymous):

\[\tau(n) = \sum_{d|n}^{} 1\]

OpenStudy (kc_kennylau):

ic

OpenStudy (kc_kennylau):

so the question becomes \(\Large\left(\sum\limits_{d|n}\sum\limits_{k|d}1\right)^2=\sum\limits_{d|n}\left(\sum\limits_{k|d}1\right)^3\)?

OpenStudy (anonymous):

i think it should be like this: \[(\sum_{d1|n}^{} \sum_{d2|n}^{} 1 )^2 = \sum_{d1|n}^{}(\sum_{d2|n}^{}1)^3\]

OpenStudy (kc_kennylau):

but d1 is the subject of the second Sigma...

OpenStudy (anonymous):

d1 is the divisors of the outer sum, d2 is the divisors for the sum of 1 so you sum all d2 first then you pick out d1 of d2

ganeshie8 (ganeshie8):

yes, d2 < d1 i think another approach is to mess with \(\tau\)and \(\sigma\) directly :- say, \(n = p_1^{k_1}p_2^{k_2}... p_r^{k_r}\) is the prime factorization of \(n\)

ganeshie8 (ganeshie8):

since, \(\tau\) is a multiplicative funciton, SUM of of \(\tau\) 's is also a multiplicative function. so it is sufficient to prove when \(n = p^k\)

OpenStudy (kc_kennylau):

but the second Sigma is not referring to the number of divisors of \(n\), but the number of divisors of \(d\)

OpenStudy (kc_kennylau):

discovery: \(\Large n=10\rightarrow(1+2+2+4)^2=1^3+2^3+2^3+4^3\) \(\Large n=15\rightarrow(1+2+2+4)^2=1^3+2^3+2^3+4^3\) Both has the same numbers... (this may seem dumb?)

ganeshie8 (ganeshie8):

interesting, but 4 is not a divisor of 10 rihgt ? the proof shrinks to proving for a single prime power \(n = p^k\) then, +ve divisors of \(n\) : \(d = {p^0, p^1, p^2, .... p^k}\) should be easy to prove

OpenStudy (kc_kennylau):

4 is not a divisor of 10, but 4 is the number of divisors of 10

ganeshie8 (ganeshie8):

gotcha :)

ganeshie8 (ganeshie8):

\(d ~~~~= {p^0, p^1, p^2, .... p^k} \) \(\tau(d) = 1, 2, 3, 4, .... k+1\)

OpenStudy (kc_kennylau):

p is prime?

ganeshie8 (ganeshie8):

\(\sum \limits_{d|n}^{} \tau(d) = 1+2+3+....+k+1 \)

OpenStudy (anonymous):

just wonder why is it sufficient to say its sufficient to prove when n = p^k yes its multiplicative why does that imply that you can prove it using prime factorizations

ganeshie8 (ganeshie8):

yup p powers are the prime factorization

OpenStudy (kc_kennylau):

but it's not prime factorization, it's factorization

ganeshie8 (ganeshie8):

cuz, \(\tau (p^kp^lp^m) =\tau (p^k) \tau (p^l)\tau (p^m)\) if u prove for one prime power, other combinations will follow.

OpenStudy (anonymous):

oh alright i get it i see what your doing now

ganeshie8 (ganeshie8):

also, \(\tau\) is multiplication implies that any sum of \(\tau\) is also multiplicative so it all reduces to proving single prime power factorization

ganeshie8 (ganeshie8):

*multiplicative

OpenStudy (kc_kennylau):

Factors of \(p^k\): 1, \(p^1\),\(p^2\), ..., \(p^k\) Taus of factors of \(p^k\): 1, 2, 3, ..., k+1

OpenStudy (anonymous):

lets see so if i prove over prime powers then \[(\sum_{d|n}^{}\tau(d))^2 = (\sum_{d|p^k}^{}\tau(p^k))^2 = \sum_{i =1}^{k} (i+1) = \sum_{i =1}^{k+1} (i) = (\frac{ (k+1)(k+2) }{ 2 })^2\]

OpenStudy (kc_kennylau):

So that's the same as proving \(\Large\left(\sum\limits_{n=1}^{k+1}n\right)^2=\sum\limits_{n=1}^{k+1}n^3\)

OpenStudy (kc_kennylau):

which is common sense lol.

ganeshie8 (ganeshie8):

looks good to me

OpenStudy (anonymous):

alright thanks for your help everyone much appreciated :D

OpenStudy (kc_kennylau):

i will obtain not any credit :) May all credit go to @ganeshie8

ganeshie8 (ganeshie8):

very good problem ! thanks for sharing wid us :)

OpenStudy (anonymous):

haha your welcome xD